The equation of the circle with center (1,2) and tangent to x+y=1.

Thanks for the big help.. Kindly put the solution..

Answer by Alan3354(21581)

The equation of the circle with center (1,2) and tangent to
x+y=1 --> y = -x + 1 (slope m = -1)
The radius will be thru the point (1,2) and perpendicular to the line x+y=1, so the slope of that radial line will be +1
y-2 = x-1
y = x+1 is the eqn of the line thru the center and perpendicular to x+y=1.
Find the intersection, which is the tangent point.
x+y=1
x-y=-1
-----
x = 0, y = 1 --> the point (0,1)
The radius is the distance from (0,1) to (1,2)
r = sqrt(2)
------------
(x-1)^2 + (y-2)^2 = 2 is the circle.